# SPDFusion: An Infrared and Visible Image Fusion Network Based on a Non-Euclidean Representation of Riemannian Manifolds

### v2

@@ Abstract
Euclidean representation learning methods have achieved commendable results in image fusion tasks, which can be attributed to their clear advantages in handling with linear space. However, data collected from a realistic scene usually have a non-Euclidean structure, where Euclidean metric might be limited in representing the true data relationships, degrading fusion performance. To address this issue, a novel SPD (symmetric positive definite) manifold learning framework is proposed for multi-modal image fusion, named SPDFusion, which extends the image fusion approach from the Euclidean space to the SPD manifolds. Specifically, we encode images according to the Riemannian geometry to exploit their intrinsic statistical correlations, thereby aligning with human visual perception. Actually, the SPD matrix underpins our network learning, with a cross-modal fusion strategy employed to harness modality-specific dependencies and augment complementary information. Subsequently, an attention module is designed to process the learned weight matrix, facilitating the weighting of spatial global correlation semantics via SPD matrix multiplication. Based on this, we design an end-to-end fusion network based on cross-modal manifold learning. Extensive experiments on public datasets demonstrate that our framework exhibits superior performance compared to the current state-of-the-art methods.
Index Terms:
image fusion, cross-modal manifold learning, statistical correlation, SPD manifold.
@@ 1 Introduction
With the development of multi-modal sensor technology, image fusion is playing an increasingly important role in the field of computer vision. This task integrates image data from different imaging sensors to obtain richer and more accurate information about the scene. Multi-modal image fusion has many applications in various practical fields such as medical imaging [1][2], remote sensing monitoring [3][4], semantic segmentation [5], and object tracking [6], etc.
@@
The introduction of many traditional methods such as multi-scale transformation [7] and low-rank representation [8][9] has significantly contributed to the task of image fusion. However, these methods typically rely on independent representations of the original features or some other forms of feature combination, such as direct use of pixel values, application of local descriptors, or the construction of histograms, which may not express the interrelationships between features effectively.
@@
Compared to the traditional image processing techniques, deep learning architectures [10] [11] [12] have shown significant effectiveness in the field of image fusion thanks to their superior ability in extracting high-level feature, learning of nonlinear transformations, and in complex pattern recognition. In the current research, methods based on Convolutional Neural Networks (CNN) [13] have shown to be effective in the field of image fusion. These methods adopt an end-to-end learning approach [14] that trains the network directly from pairs of input images to produce the fused result, capturing image characteristics at different levels from low-level edges and textures to higher-level semantic content. However, CNN extracts local features through convolution operations in the local receptive field, and this often leads to the neglect of long-range dependencies in the image, which is disadvantageous for achieving high-quality image fusion. For instance, in tasks such as depth-of-field extension or multi-focus image fusion [15] [16], maintaining relations between foreground and background usually requires considering contextual information from the entire image scene.
Figure 1: A new modeling approach based on the representation of Riemannian manifold. We divide the images from two sources of data into patches, allowing the network to focus on local details and global context information.Where the cross-modal image pairs are considered to be on the manifold $\mathcal{S}_{1}$, while the manifold $\mathcal{S}_{2}$ where the patches reside can be regarded as a manifold structure that is homeomorphic to $\mathcal{S}_{1}$. $\mathcal{M}$ denotes the SPD Riemannian manifold, $Sym_{++}^{d}$ represents the symmetric positive definite manifold, and $T_{I}Sym_{++}^{d}$ is the tangent space of the SPD manifold at the $d\times d$ identity matrix $I$.
@@
In 2021, the Vision Transformer (ViT) [17] appeared as a revolutionary architecture, shifting the deep learning paradigm from traditional convolution operations to processing images by dividing them into fixed-size tokens, which can capture long-range dependencies between image segments. Subsequent applications, especially in image fusion tasks, such as TransMEF [18] and SwinFusion [19], capitalized on this method’s powerful capabilities. Despite the significant progress, these models overly focus on the self-attention mechanism itself, sacrificing cross-modal interaction processes. Yet, in fusion tasks, such associative information is crucial[20] [21].
In computer vision, data often contains rich intrinsic structures and hidden associations, especially in multi-modal fusion tasks. Samples within a modality are highly correlated in the feature space, while the relationship between samples across different modalities is weaker.
@@
To tap the implicit interactions and dependencies of data from different modalities, bearing in mind the intrinsic non-Euclidean manifold structure of images, we propose an infrared and visible light image fusion algorithm based on the SPD manifold. This model applies Riemannian manifold [12][22] to the image fusion task for the first time, mining the raw pixel characteristics and their statistical relationships to guide the fusion results. This method not only preserves the inherent geometric properties of image data on the Riemannian manifold but also facilitates the interaction of cross-modal image correlation information. On one hand, we design an SPD manifold attention module (SPDAM), which, through the introduction of a mixed attention mechanism, strengthens the statistical correlation within each modality and across the different modalities. On the other hand, geometric computations on Riemannian manifolds enable a more refined capture and utilization of the spatial nonlinear characteristics, thereby providing a better fit to the intricate geometric structure inherent in image data. Based on this, we model the covariance among image patches, effectively achieving the learning of global correlation relationships, and the main idea of the proposed method is shown in Fig. 1. Our main contributions are as follows:
@@
1. For the first time, the Riemannian manifold is introduced into the image fusion task. The proposed method exhibits unique advantages due to its adaptation to the non-Euclidean distribution of image data and the consideration of its intrinsic geometric structure.
@@
2. A novel manifold attention module has been designed, which reflects the global correlation between features through the covariance of pixel point representations. As the manifold network continuously updates, it further strengthens semantic associations, effectively guiding the fusion task.
@@
3. A novel cross-modal fusion strategy has been introduced, which integrates features from different modalities on the SPD manifold, enhancing the inter-modality correlations and improving the fusion effectiveness.
@@
4. Experimental results on public benchmark datasets indicate that our fusion network exhibits superior fusion performance compared to existing fusion methods.
2 Related works
In this section, we explore neural networks based on covariance matrices as well as SPD neural networks constructed upon them, and also elaborate on network modules based on spatial attention.
@@ 2.1 Covariance Matrix
The covariance matrix is an important tool for describing the statistical linear dependence between variables in a dataset [23]. It offers a comprehensive perspective on multivariate data in terms of of second-order moment statistics and provides a mathematically rigorous method for quantitative modeling of the interrelationships between variables.
The potential of the covariance matrix has been effectively utilized in many branches of statistics. However, its application in pattern recognition is relatively rare. It has been proven that covariance matrices are effective in various scenarios. For example, in image set classification [24] [25] [22], the covariance matrix can represent the statistical characteristics of a set of images and characterize the image set as a distribution of feature points. This property of covariance can be used to perform more complex tasks, such as recognizing facial expressions or motion patterns. In medical imaging [26], the covariance matrix is used to capture morphological changes in various tissues or structures, and by studying these changes, disease characteristics can be revealed, helping doctors make more accurate diagnoses.
However, Riemannian networks based on covariance matrices are mainly applied to classification problems, and the features extracted by them often do not fully meet the specific requirements of image fusion tasks. Additionally, the effect of regional correlations between inter-modal data on the fusion outcome has not been fully understood and studied, particularly in tasks involving multi-modal data fusion, such as analyzing experimental data, multi-spectral satellite images, or multi-modal medical images. These correlations can reveal the intrinsic connections between the data of different modalities, such as the scene features captured by different sensors, or help distinguish unrelated information. This is crucial for building accurate fusion models and for understanding the performance of the data after fusion.
Figure 2: The workflow of our SPDFusion: (a) The main framework of our fusion network is defined. (b) Details of manifold learning. (c) The Riemannian attention weighting operation, which is tailored specifically for our fusion task.
@@ 2.2 The Basic Layer of SPD Neural Network
Since traditional machine learning and classification methods generally assume the Euclidean space, these methods are not suitable for dealing with the nonlinear structure of covariance matrices, as they fail to take into account the intrinsic geometric characteristics of data points on the manifold. Fortunately, researchers have proposed an innovative neural network architecture: SPDNet [27], which is specifically designed to adapt to the unique features of SPD matrices. In its overall structure, it is similar to traditional Euclidean neural networks, including transformations, nonlinear activations, and the final classification stage. The difference lies in the fact that the SPDNet architecture includes specialized layers to ensure the outputs remain on the manifold, utilizing operations such as Riemannian metric, exponential and logarithmic mappings, as well as geodesic distances compatible with manifold geometry [28]. The basic layers can be divided into three categories:
@@
$X_{k}=f_{b}(X_{k-1};W_{k})=W_{k}X_{k-1}W_{k}^{T}$ (1)
Where $X_{k}$ is the newly mapped SPD matrix, $X_{k-1}$ is the SPD matrix input at the $k$-th layer, belonging to the manifold of symmetric positive definite matrices, $W_{k}$ is the transformation matrix, which should be row full-rank, and if its columns are orthogonal, $W_{k}$ belongs to the compact Stiefel manifold $St(d_{k},d_{k-1})$, where $d_{k}$ is the dimension of the mapped SPD matrix after the transformation, typically smaller than $d_{k-1}$, to achieve dimensionality reduction.
@@
$A(i,i)=\begin{cases}\sum_{k-1}(i,i)&\text{if }\sum_{k-1}(i,i)>\epsilon\\ \epsilon&\text{if }\sum_{k-1}(i,i)\leq\epsilon\end{cases}$ (3)
@@
The key to the LogEig layer is that it offers a means to effectively ”straightening” the curved manifold space into a flat Euclidean space while preserving the SPD property (positive-definiteness) by performing logarithmic operations on the eigenvalues obtained from an EIG (eigenvalue decomposition) of the matrix. This makes linear operations feasible in the Euclidean space.
$X_{k}=f_{l}^{(k)}\left(X_{k-1}\right)=\log\left(X_{k-1}\right)=U_{k-1}\log\left(\Sigma_{k-1}\right)U_{k-1}^{T}$ (4)
@@
For a SPD matrix $X_{k-1}$, one first performs the eigenvalue decomposition $X_{k-1}=U_{k-1}\Sigma_{k-1}U_{k-1}^{T}$, where $U_{k-1}$ is an orthogonal matrix comprised of eigenvectors, and $\Sigma_{k-1}$ is a diagonal matrix of corresponding eigenvalues. Following this, the logarithm of the diagonal elements of $\Sigma_{k-1}$ is computed to obtain the diagonal matrix of log-eigenvalues $\log(\Sigma_{k-1})$. Finally, the SPD matrix $X_{k}$ is reconstructed through $U_{k-1}\log(\Sigma_{k-1})U_{k-1}^{T}$, thus $X_{k}$ represents $X_{k-1}$ in the log domain.
2.3 Deep Learning Tasks Based on Spatial Attention
The concept of attention mechanisms is derived from the basic characteristics of the human perception system. It is well known that human vision does not uniformly process the entire field of view; instead, it prioritizes those parts that are most informative. This selective focus allows us to efficiently manage our limited cognitive resources to better understand our environment.
In the fields of deep learning and computer vision, attention mechanisms attempt to mimic this selective visual focus. The traditional Convolutional Neural Networks (CNNs) [29] [30] [31] typically treat all parts of the input equally, which may not be the most efficient method when dealing with complex or cluttered data. With the development of large-scale classification tasks and other vision tasks, researchers began to explore how to make CNNs focus more on important local information in the image, similar to human vision.
In image processing tasks, spatial attention and channel attention are two common types of attention. Spatial attention focuses on the importance of specific areas within an image, while channel attention focuses on the importance of feature channels.
@@ 2.3 Deep Learning Tasks Based on Spatial Attention
Hu et al. proposed a structural unit called SEBlock (Squeeze-and-Excitation block) [32] based on the inter-relationships between channels. It dynamically calibrates channel feature responses by explicitly modeling the dependencies between channels. However, the SEBlock’s approach is limited in that it uses global average pooling to calculate channel attention, which may not be the optimal feature for inferring fine-grained channel attention, and it does not consider the equally important spatial attention, thus failing to achieve semantic-level perception of the context of objects in a scene and their spatial relationships.
@@
CBAM (Convolutional Block Attention Module) [33] introduces spatial attention on top of this. It is a lightweight attention module that refines the input feature maps by sequentially applying channel and spatial attention mappings. Unlike channel attention, spatial attention focuses on the ”where” question in the input feature maps, identifying which areas of the image parts are information-rich, in order to emphasize or suppress these areas. This is usually achieved by processing each spatial location of the feature maps to generate a spatial attention map that indicates the importance of each location.
In traditional Euclidean space attention mechanisms, the attention weights are calculated through dot-product operations between query and key vectors. This method shows limitations when dealing with data that contains complex manifold structures. To address this, our proposed manifold attention expands the traditional Euclidean attention to data representations based on symmetric positive definite manifolds. the key innovation is to learn latent representational associations between two modalities through covariance modeling, thereby establishing a new paradigm for representing global features statistically. More critically, this framework is optimized within a Riemannian manifold space, allowing the algorithm to handle non-linear structures in a more natural manner, and preserving the geometric relationships and topological properties of the data.

### v3

@@ Abstract
Euclidean representation learning methods have achieved promising results in image fusion tasks, which can be attributed to their clear advantages in handling with linear space. However, data collected from a realistic scene usually has a non-Euclidean structure, evaluating the consistency of latent representations from paired views using Euclidean distance raises challenges. To address this issue, a novel SPD (symmetric positive definite) manifold learning is proposed for multi-modal image fusion, named SMLNet, which extends the image fusion approach from the Euclidean space to the SPD manifolds. Specifically, we encode images according to the Riemannian geometry to exploit their intrinsic statistical correlations, thereby aligning with human visual perception. The SPD matrix fundamentally underpins our network’s learning process. Building upon this mathematical foundation, we employ a cross-modal fusion strategy to exploit modality-specific dependencies and augment complementary information. To capture semantic similarity in images’ intrinsic space, we further develop an attention module that meticulously processes the cross-modal semantic affinity matrix. Based on this, we design an end-to-end fusion network based on cross-modal manifold learning. Extensive experiments on public datasets demonstrate that our framework exhibits superior performance compared to the current state-of-the-art methods. Our code will be publicly available at https://github.com/Shaoyun2023.
@@ 1 Introduction
With the development of multi-modal sensor technology, image fusion is playing an increasingly important role in the field of computer vision. This task integrates image data from different imaging sensors to obtain richer and more accurate information about the scene. Multi-modal image fusion has many applications in various practical fields such as medical imaging  (Xu and Ma, 2021; Li et al., 2021b), remote sensing monitoring  (Ma et al., 2020; Zhang et al., 2020a), semantic segmentation  (Tang et al., 2022a), and object tracking  (Zhu et al., 2024), etc.
@@
The introduction of many traditional methods such as multi-scale transformation  (Chen et al., 2020) and low-rank representation  (Zhang et al., 2021b; Li et al., 2020, 2023) has significantly contributed to the image fusion task. However, these methods typically depend on isolated feature representations or basic combinations, such as raw pixel values, local descriptors, or histograms, none of which adequately model global feature relationships. Additionally, they neglect the interactions among local features. To address this, our method introduces the covariance matrix as a mathematical descriptor to quantify statistical dependencies among feature blocks, identifying positive and negative correlations of features. This is essential for uncovering the latent dependencies between global and local features in an image.
@@
Compared to the traditional image processing techniques, deep learning architectures (Zhang et al., 2021a; Zhang and Demiris, 2023; Wang et al., 2023c; Liu et al., 2017) have shown significant effectiveness in the field of image fusion due to their superior ability in extracting high-level image features, and learning of nonlinear transformations. In the current research, methods based on Convolutional Neural Networks (CNNs)  (Zhang et al., 2020b; Liu et al., 2017) have shown to be effective in the field of image fusion. These methods adopt an end-to-end learning paradigm  (Li et al., 2021a; Deng et al., 2024) that trains the network directly from pairs of input images to produce the fused result, capturing image characteristics at different levels from low-level appearance information to higher-level semantic content. However, CNN extracts image features through convolution operations in the local receptive field, and it underestimates the long-range dependencies in the image, which impedes the image fusion performance. For instance, in multi-focus image fusion  (Liu et al., 2020, 2017), maintaining relations between foreground and background usually requires considering contextual information from the entire image scene.
@@
In 2021, the Vision Transformer (ViT)  (Dosovitskiy et al., 2021) appeared as a revolutionary architecture, shifting the deep learning paradigm from traditional convolution operations to processing images by dividing them into fixed-size tokens, which can capture long-range dependencies between image segments. Subsequent applications, especially in image fusion tasks (Qu et al., 2022; Ma et al., 2022; Zhao et al., 2023a; Li and Wu, 2024; Cheng et al., 2025c), capitalized on this method’s powerful capabilities. Although significant progress has been made, current vision attention models extremely rely on Euclidean algebraic operations, causing the output of attention to neglect the geometric dependencies within high-dimensional data.
In computer vision, data often exhibits rich intrinsic structures and hidden associations, particularly in multi-modal fusion tasks where samples within a single modality show strong feature-space correlations, while cross-modal relationships remain sparse and underutilized. This gap is effectively bridged by the SPD manifold, composed of symmetric positive definite matrices, which inherently preserves second-order statistical characteristics through its geometric structure. By doing so, it becomes ideal for modeling cross-modal dependencies, where covariance directly quantifies shared variations (e.g., thermal-texture alignment), enabling precise statistical interaction modeling even when inter-modal relationships are weak.
To effectively model intrinsic cross-modal statistical correlations, our approach leverages the geometric properties of SPD manifolds to represent image features, overcoming the limitations of traditional Euclidean methods. By constructing composite covariance matrices that decompose relationships into self-correlation and cross-correlation terms, we explicitly quantify implicit intra-modal and inter-modal dependencies while preserving the manifold’s geometric constraints (symmetry, positive definiteness). Unlike Euclidean approaches, this SPD-based paradigm enables statistically rigorous fusion through intrinsic manifold operations, ensuring that the fused feature representations not only capture nonlinear interactions but also align with the true distribution of multi-modal data.
Figure 1: A new modeling approach based on the representation of Riemannian manifold. We divide the images from two sources of data into patches, allowing the network to focus on local details and global context information. Where the cross-modal image pairs are considered to be on the manifold $\mathcal{S}_{1}$, while the manifold $\mathcal{S}_{2}$ where the patches reside can be regarded as a manifold structure that is homeomorphic to $\mathcal{S}_{1}$. $\mathcal{M}$ denotes the SPD Riemannian manifold, $Sym_{++}^{d}$ represents the symmetric positive definite manifold, and $T_{I}Sym_{++}^{d}$ is the tangent space of the SPD manifold at the $d\times d$ identity matrix II.
@@
Based on this, we propose an infrared and visible image fusion algorithm based on the SPD manifold. This model pioneers the application of Riemannian manifold networks  (Thanwerdas and Pennec, 2023; Brooks et al., 2019; Nguyen, 2021; Huang and Van Gool, 2017) to multi-modal image fusion tasks, mining the raw pixel characteristics and their statistical relationships to guide the fusion process. This method not only preserves the inherent geometric properties of image data on the Riemannian manifold but also facilitates the interaction of cross-modal information. On one hand, we design an SPD manifold attention module (SPDAM), which, through the introduction of a mixed attention mechanism, strengthens the statistical correlation within each modality and across different modalities. On the other hand, geometric computations on Riemannian manifolds enable a more refined capture and utilization of the spatial nonlinear characteristics, thereby providing a better fit to the intricate geometric structure inherent in image data. Based on this, we model the covariance among image patches, effectively achieving the learning of global correlation relationships, and the main idea of the proposed method is shown in Fig. 1. Our main contributions are as follows:
The contributions of this work can be summarized as follows:
@@
  • For the first time, the Riemannian manifold network is introduced into the image fusion task. The proposed method exhibits unique advantages due to its adaptation to the non-Euclidean distribution of image data and the consideration of its intrinsic geometric structure.
@@
  • A new manifold attention module is designed. Through global correlation representation learning, it strengthens the semantic correlation between features and effectively guides the fusion task.
@@
  • A novel cross-modal fusion strategy has been introduced, which integrates features on the SPD manifold to enhance intrinsic semantic coherence across modalities.
@@
  • Experimental results on public benchmark datasets indicate that our fusion network exhibits superior fusion performance compared to existing fusion methods.
2 Related Work
@@ 2.1 Covariance Matrix
The covariance matrix is an important tool for describing the statistical linear dependence between variables in a dataset  (Wang et al., 2012). It offers a comprehensive perspective on multivariate data in terms of of second-order moment statistics and provides a mathematically rigorous method for quantitative modeling of the interrelationships between variables.
The potential of the covariance matrix has been effectively utilized in many branches of statistics. However, its application in pattern recognition is relatively underestimated. It has been proven that covariance matrices are effective in various scenarios. For example, in image set classification  (Wang et al., 2022a)(Wang et al., 2022b), the covariance matrix can represent the statistical characteristics of a set of images and characterize the image set as a distribution of feature points. This property of covariance can be used to perform more complex tasks, such as recognizing facial expressions or motion patterns. In medical imaging  (Chen et al., 2023), the covariance matrix is used to capture morphological changes in various tissues or structures, and by studying these changes, disease characteristics can be revealed, helping doctors make more accurate diagnoses.
However, Riemannian networks using covariance matrices are primarily designed for classification tasks, and their extracted features often fall short of meeting the specific demands of image fusion. Moreover, the impact of inter-modal regional correlations on fusion results remains understudied, especially in multi-modal scenarios like multi-spectral satellite imagery and medical image fusion. These correlations can uncover relationships between different modalities, such as scene features from diverse sensors, or help filter irrelevant information.
Considering the advantages of covariance modeling, Our SMLNet innovatively employs a composite covariance matrix architecture to mathematically model both intra-modal feature distributions (e.g., thermal radiation in infrared images and texture details in visible images) and inter-modal joint distribution patterns. This dual statistical modeling mechanism not only completely preserves the unique feature representations of each modality, but more importantly establishes precise statistical dependencies between cross-modal features, thereby achieving efficient suppression of redundant information while enhancing complementary features during the fusion process.
@@ 2.2 The Basic Layer of SPD Neural Network
Since traditional machine learning and classification methods operate primarily in Euclidean space, they are ill-suited for processing SPD matrices. These methods fail to account for the intrinsic Riemannian manifold structure of SPD matrices and may distort their geometric properties. Fortunately, researchers have proposed an innovative neural network architecture: SPDNet  (Huang and Van Gool, 2017), which is specifically designed to adapt to the unique features of SPD matrices. In its overall structure, it is similar to traditional Euclidean neural networks, including transformations, nonlinear activations, and the final classification stage. The difference lies in the fact that the SPDNet architecture includes specialized layers to ensure the outputs remain on the manifold, utilizing operations such as Riemannian metric, exponential and logarithmic mappings, as well as geodesic distances compatible with manifold geometry  (Brooks et al., 2019). The basic layers can be divided into three categories:
@@
$X_{k}=f_{b}(X_{k-1};W_{k})=W_{k}X_{k-1}W_{k}^{T},$ (1)
where $X_{k}$ is the newly mapped SPD matrix and $X_{k-1}$ is the input SPD matrix to the $k$-th layer. $W_{k}$ is the transformation matrix, which should be row full-rank, and if its columns are orthogonal, $W_{k}$ belongs to the compact Stiefel manifold $St(d_{k},d_{k-1})$, where $d_{k}$ is the dimension of the mapped SPD matrix after the transformation, typically smaller than $d_{k-1}$, to achieve dimensionality reduction.
@@
[…]
@@
The key to the LogEig layer is that it offers a means to effectively “straightening” the curved manifold space into a flat Euclidean space while preserving the SPD property (positive-definiteness) by performing logarithmic operations on the eigenvalues obtained from an EIG (eigenvalue decomposition) of the matrix. This makes linear operations feasible in the Euclidean space:
$X_{k}=f_{l}^{(k)}\left(X_{k-1}\right)=\log\left(X_{k-1}\right)=U_{k-1}\log\left(\Sigma_{k-1}\right)U_{k-1}^{T},$ (4)
@@
where for a SPD matrix $X_{k-1}$, one first performs the eigenvalue decomposition $X_{k-1}=U_{k-1}\Sigma_{k-1}U_{k-1}^{T}$, where $U_{k-1}$ is an orthogonal matrix comprised of eigenvectors, and $\Sigma_{k-1}$ is a diagonal matrix of corresponding eigenvalues. Following this, the logarithm of the diagonal elements of $\Sigma_{k-1}$ is computed to obtain the diagonal matrix of log-eigenvalues $\log(\Sigma_{k-1})$. Finally, the SPD matrix $X_{k}$ is reconstructed through $U_{k-1}\log(\Sigma_{k-1})U_{k-1}^{T}$, thus $X_{k}$ represents $X_{k-1}$ in the log domain.
Figure 2: The workflow of our SMLNet: (a) The main framework of our fusion network is defined. (b) Details of manifold learning. (c) The Riemannian attention weighting operation, which is tailored specifically for fusion task. Our approach begins by leveraging the non-Euclidean structure of the data and modeling it in manifold space. The manifold network aggregates cross-modal information while the SPD Attention Module (SPDAM) dynamically allocates attention weights to adaptively align cross-modal image features. The final output is generated through a weighted fusion of the two modal convolutional components.
2.3 Deep Learning Tasks Based on Attention Mechanism
In deep learning and computer vision, attention mechanisms aim to emulate the human ability to focus selectively on relevant visual information. While traditional CNNs process all input regions uniformly, such an approach can be suboptimal for complex or cluttered data, see  (Wang et al., 2018; Zhang et al., 2019). With the development of large-scale classification tasks and other vision tasks, researchers began to explore how to make CNNs focus more on important local information in the image to learn discriminative feature representations. This pursuit has led to the development of two principal attention paradigms: spatial attention and channel attention, which are widely used in image processing tasks. Specifically, spatial attention focuses on identifying the importance of specific spatial regions within an image, whereas channel attention operates by emphasizing critical feature channels.
@@ 2.3 Deep Learning Tasks Based on Attention Mechanism
Hu et al. proposed a structural unit called SEBlock (Squeeze-and-Excitation block)  (Hu et al., 2018) based on the inter-relationships between channels. It dynamically calibrates channel feature responses by explicitly modeling the dependencies between channels. However, the SEBlock’s approach is limited in that it uses global average pooling to calculate channel attention, which may not be suitable for inferring fine-grained channel attention, and it does not consider the equally important spatial attention, thus failing to achieve semantic-level perception of the context of objects in a scene.
@@
CBAM (Convolutional Block Attention Module)  (Woo et al., 2018) introduces spatial attention on top of this. It is a lightweight attention module that refines the input feature maps by sequentially applying channel and spatial attention mappings. Unlike channel attention, spatial attention focuses on the “where” question in the input feature maps, identifying which areas of the image parts are information-rich, in order to emphasize or suppress these areas. This is usually achieved by processing each spatial location of the feature maps to generate a spatial attention map that indicates the importance of each location.
Traditional Euclidean attention mechanisms, fundamentally constrained by their reliance on linear dot-product operations, fail to capture intrinsic statistical correlations in data with complex manifold structures. We overcome this limitation by generalizing attention computation to SPD manifolds, explicitly modeling higher-order statistical dependencies via covariance interactions across modalities. Through Riemannian manifold optimization, SMLNet ensures preservation of multi-modal geometric priors while capturing nonlinear feature covariances, thereby constructing geometry-aware cross-modal representations.